# DriveGPT: Scaling Autoregressive Behavior Models for Driving

### v1

@@ Abstract
We present DriveGPT, a scalable behavior model for autonomous driving. We model driving as a sequential decision making task, and learn a transformer model to predict future agent states as tokens in an autoregressive fashion. We scale up our model parameters and training data by multiple orders of magnitude, enabling us to explore the scaling properties in terms of dataset size, model parameters, and compute. We evaluate DriveGPT across different scales in a planning task, through both quantitative metrics and qualitative examples including closed-loop driving in complex real-world scenarios. In a separate prediction task, DriveGPT outperforms a state-of-the-art baseline and exhibits improved performance by pretraining on a large-scale dataset, further validating the benefits of data scaling.
@@ 1 Introduction
In the context of autonomous driving, transformer-based models have demonstrated great success in behavior modeling [7, 23, 22]. Behavior modeling predicts future actions of traffic agents to support critical tasks such as planning and motion prediction. Unlike traditional machine learning approaches that may rely on handcrafted features or specific rules, transformers are effective at learning the underlying spatiotemporal dependencies of agent behavior from large-scale multimodal data.
Scaling up model and dataset sizes has been critical for recent advances in sequential modeling for text prediction [12], yet it remains an open question whether these scaling trends hold for the task of behavior modeling. Existing work is often limited by the training data or the model size, as summarized in Table 1.
@@
In our work, we present a comprehensive study of scaling up data sizes and model parameters in the context of autonomous driving. Specifically, we train a transformer-based autoregressive behavior model on over 100 million high-quality human driving examples, $\sim$50 times more than existing open-source datasets, and scale the model over 1B parameters, outsizing existing published behavior models.
@@
    We determine empirical driving scaling laws for an autoregressive behavior model in terms of model parameters, data size, and compute. We validate the value of more training data through data scaling that spans two orders of magnitude and compute scaling that spans five orders of magnitude. Our model scaling experiments demonstrate better scalability of autoregressive decoders compared to one-shot decoders.
@@
    We quantitatively and qualitatively compare models from our scaling experiments to validate their effectiveness in real-world driving scenarios.
@@ 2.1 Behavior Modeling
Behavior modeling is a critical task in autonomous driving, which covers a broad spectrum of tasks including planning, prediction, and simulation. Taking multimodal inputs including agent history states and map information, behavior models predict the future states of these traffic agents, by reasoning agent dynamics [2, 24], interactions [25, 10], and driving environments [13, 15].
Among all learning-based models, transformers have gained popularity due to their capability to fuse multimodal inputs as encoders [18, 30, 32, 9] and modeling long-range temporal relationships as decoders [22, 23]. Despite the success of transformers in behavior modeling, existing literature is often restricted by the size of model parameters due to limited training data, which fail to capture the full scaling potential of transformer-based models. In our work, we scale up our transformer models to include billions of parameters, by training on a large-scale dataset including more than 100M driving demonstrations, and validate the scalability of transformer-based models in the context of autonomous driving.
@@ 2.2 Large Transformer Models
Recent work has studied the scalability of behavior models in the context of motion prediction [22], planning [26], and simulation [8], yet these studies are either constrained by limited data size (up to a couple of million training examples), or focused on a few orders of magnitude in terms of data and model scaling, limiting the potential to draw statistically significant conclusions over a large scaling spectrum.
In this paper, we aim to study the scaling properties in terms of model parameters, data samples, and compute, across a much larger spectrum compared to prior work. More specifically, we target an autoregressive decoder architecture that has been proven to be both scalable as in the LLM literature and effective in generating accurate trajectories for traffic agents [22].
Beyond autonomous driving, there exists a few relevant literature on building large transformer models for robotic tasks [19, 27], which share a similar transformer architecture as our work. Despite their successes, robotics models often share different input features and dynamic models, and operate in different environments, making them difficult to apply directly to driving tasks.
@@ 3 Behavior Model
We use a standard encoder-decoder architecture as our behavior model, as shown in Fig. 2. We use transformer-based models as our encoder and decoder backbones due to their scalability in sequential modeling tasks.
@@ 3.3 LLM-Style Trajectory Decoder
where $a_{t}$ is the predicted Verlet action, and $(s_{t}-s_{t-1})$ is the estimated velocity. This representation helps predict smooth trajectories using a small set of action candidates.
@@ 3.4 Training
We use a single cross-entropy classification loss over the action space, where the target action is selected as the one that is closest to the ground truth future trajectories. We refer the readers to the Appendix for more details.
@@ 4 Scaling Experiments
The goal of our scaling experiments is to determine the effect of model and data size on behavior prediction performance. Quantifying scaling laws similar to those seen in LLMs [12] can help prioritize the value of data and compute for future research directions in behavior modeling. We focus our effort on exploring the next frontier of model and data size — over an order of magnitude beyond previously reported work.
Large-scale driving dataset
From millions of miles of high-quality real-world human driving demonstrations, we curated a small subset of 120M segments for an internal research dataset. The dataset is balanced to cover diverse geographic regions from multiple cities and countries, and a variety of driving scenarios, such as passing a double-parked vehicle, unprotected left turns, construction zones, etc. We extracted map information, target agent states, and nearby agent states into vectorized representation, as customary in behavior modeling literature [4].
@@ 4.1 Data Scaling
Figure 3: Performance increases with dataset size across a range of model parameters, indicating that data is a limiting factor. Both axes are on a log scale. An exponential fit was applied to all data points except for the 1.5M curve, resulting in the following relationship: $\log(L)=-0.101\log(D)+2.656$ with an $R^{2}$ value of 0.9902, where $L$ is the validation loss and $D$ is the number of unique training samples.
Data scaling results are summarized in Fig. 3. The smallest dataset of 2.2M samples mimics the size of Waymo Open Motion Dataset (WOMD) [3], a large open-source dataset for behavior modeling ($\sim$44k scenarios with multiple target agents per scenario). Additionally, we select a few subsets of our internal research dataset to study data scaling across different orders of dataset sizes. Our experiments use $\sim$50x more data than WOMD, exploring a new region of the design space.
The results indicate that as the model is trained on more unique data samples, the performance improves, regardless of model size. Extrapolating from the scaling law in Fig. 3, to improve the best loss by another $10\%$, we would need to include $200$M more training examples. A $20\%$ improvement would require about $900$M more examples. As a result, we find that data remains to be the bottleneck for further improving driving performance.
@@ 4.2 Model Scaling
Figure 4: Validation loss saturates with increasing model size across all datasets — in contrast to LLM scaling literature.
We now study model sizes across three orders of magnitude (1.5M to 1.4B parameters). We increase model size by increasing the hidden dimensions of transformers for simplicity. We notice that modifying other parameters such as number of attention heads and hidden dimensions per head do not lead to noticeable changes in the results.
Training larger models is sensitive to learning rates, as observed in other scaling studies [12, 6]. For each model size, we run multiple experiments at different learning rates to select the one with the optimal performance. We defer more detail in the Appendix.
Results in Fig. 4 show that for a fixed dataset size, increasing model parameters has a mixed effect on the validation loss. The losses improve up to 26M parameters, before hitting a plateau up to 163M parameters. While these results contrast with the LLM scaling literature [12], which shows significant benefits from model scaling, we notice that actual driving metrics continue to improve as model scales, as shown in Sec. 4.4 and Sec. 5.1.2. Further investigation is required to determine the gap between loss and metric measures, and whether there exist fundamental differences in the behavior prediction vs text prediction task.
@@ 4.3 Compute Scaling
Next, we investigate if there is an optimal combination of model parameters and data size for a fixed compute budget. Given the large computational expense for training models at the scales we are exploring, it is important to make the best use of our data. In this study, we fixed the compute budget at different FLOP groups. For a fixed compute budget, we can allocate resources either to model parameters or data samples, keeping their product constant.
@@
Fig. 6 plots performance with different compute budgets. The trend clearly shows that a larger compute budget leads to better performance, with optimal model size increasing accordingly, as indicated by the “best” gray line. The results further reveal that data is the main bottleneck, as the smallest model size performs best at the three largest FLOP groups.
@@ 5 Planning and Prediction Experiments
5.1 Internal Evaluation: AV Planning
We measure the planning performance on a comprehensive test set through a set of standard geometric metrics including minADE (mADE), minFDE (mFDE), and miss rate (MR). Additionally, we use semantic-based metrics including offroad rate (Offroad) that measures the ratio of trajectories that leave the road and collision rate (Collision) that measures the ratio of trajectories overlapping with static agents. We normalize these metrics across experiments to highlight relative performance changes as data size and model parameters scale.
@@ 5.1.1 Data Scaling Results
We compare a DriveGPT model at 26M parameters (near-optimal from Sec. 4) trained on datasets of different sizes. The baseline dataset (2.2M) is selected to mimic the training size of a typical behavior modeling dataset such as WOMD.
The results are presented in Table 2, where we see that training on more data samples significantly improves the quality of predicted AV trajectories, in terms of critical semantics metrics in driving including offroad rate and collision rate, as well as geometric metrics. These results are consistent with Sec. 5.1.1.
@@ 5.1.2 Model Scaling Results
The results are presented in Table 3, where all metrics improve as the model size increases up to 94M parameters, with further gains in the collision metric at 163M parameters. Although validation loss stabilizes between 26M and 163M in Fig. 4, driving metrics continue to improve with increased model capacities, highlighting the potential benefits of introducing more parameters for enhanced driving performance.
@@ 5.1.3 Closed-Loop Driving
We demonstrate the effectiveness of DriveGPT as a real-time motion planner in a closed loop setting. The model takes input features from an industry-level perception system that outputs agent states and map information.
In Fig. 10, we present a challenging example in dense urban traffic where there are two double-parked vehicles blocking the path forward along with other oncoming vehicles.  DriveGPT generates smooth and safe trajectories, bypassing the blocking vehicles and moving back to the original lane afterward. More examples are presented in the supplementary material.
@@ 5.2 External Evaluation: Motion Prediction
MTR [23] 0.3292 0.6009 0.1087 0.5400 0.6655 1.3155 0.1463 0.4548 1.2454 2.6178 0.1935 0.3795
DriveGPT-WOMD 0.2875 0.5131 0.0842 0.3951 0.5716 1.0972 0.1180 0.3074 1.0553 2.2434 0.1676 0.2546
DriveGPT-Finetune 0.2866 0.5101 0.0817 0.3825 0.5659 1.0795 0.1136 0.2984 1.0404 2.2069 0.1631 0.2497
Table 4: Compared to a state-of-the-art baseline method, our model achieves better geometric metrics for vehicle prediction using an autoregressive decoder architecture, and even better metrics after pretraining on our internal research dataset.
@@
MTR [23] 0.3440 0.7193 0.0733 0.4228
DriveGPT-WOMD 0.3086 0.6346 0.0732 0.2908
DriveGPT-Finetune 0.3048 0.6253 0.0696 0.2810
@@
MTR [23] 0.7045 1.4248 0.1836 0.3653
DriveGPT-WOMD 0.6351 1.2849 0.1752 0.2667
DriveGPT-Finetune 0.6253 1.2593 0.1717 0.2339
Table 5:  DriveGPT achieves better geometric results in pedestrian and cyclist prediction. Results are averaged over 3 prediction horizons.
@@ 5.2.1 Open-Source Encoder
For our external evaluation, we use the open-source MTR [23] encoder. This encoder is similar to the one described in Sec. 3.2. We make this change to improve the reproducibility of our results and take advantage of MTR’s open-source dataloading code for WOMD. We use the same autoregressive decoder following [22], as described in Sec. 3.3.
@@ 5.2.2 Pretraining Setup
We pretrain DriveGPT by training on our internal research dataset for one epoch (as in Sec. 4). We load the pretrained checkpoint and finetune the model using the same training setup as in the MTR codebase [23], where we train the model for 30 epochs using a weighted decay learning rate scheduler.
5.2.3 Validation Results
@@
We measure model performance via a set of standard WOMD metrics, including minADE (mADE), minFDE (mFDE), miss rate (MR), and mAP. Each metric is measured on the validation set and computed over three different time horizons.
@@
We select MTR [23], a state-of-the-art motion prediction model, as our baseline and reproduce its results using its open-source code. We present two variants of our method to validate its effectiveness on the motion prediction task, including DriveGPT-WOMD that is trained on WOMD, and DriveGPT-Finetune that is pretrained on our 120M internal research dataset and finetuned on WOMD.
@@
The vehicle prediction results are summarized in Table 4, where we see that our model outperforms MTR in all geometric metrics by a large margin, especially at short horizons (2s and 5s) where we see more than 12% improvements in minADE and minFDE, as well as 20% in miss rate. While we focus more on geometric metrics that emphasize the recall of predicted trajectory samples (i.e. not missing the critical trajectory), our model shows lower mAP scores due to suboptimal probability estimates. These estimates are computed by accumulating log probabilities over a long sequence, which reveals a limitation of using an autoregressive decoder to produce accurate probability estimates, as also noted in the LLM literature [11, 5], and we defer improving the probability estimates as future work.
@@
We observe up to 4% additional gains by pretraining on our internal dataset, despite a large distribution shift between our internal dataset and the public WOMD dataset, in terms of trajectory distributions, feature noises, and differences in semantic definitions.
Beyond vehicle prediction, we report results in predicting pedestrians and cyclists in Table 5, where we see a similar level of improvements. This validates the generalizability of our method of predicting a diverse set of agent types. Again, we see extra gains by pretraining on our internal research dataset, especially for cyclists.
5.2.4 Data Scaling in Pretraining

### v2

@@ Abstract
We present DriveGPT, a scalable behavior model for autonomous driving. We model driving as a sequential decision-making task, and learn a transformer model to predict future agent states as tokens in an autoregressive fashion. We scale up our model parameters and training data by multiple orders of magnitude, enabling us to explore the scaling properties in terms of dataset size, model parameters, and compute. We evaluate DriveGPT across different scales in a planning task, through both quantitative metrics and qualitative examples, including closed-loop driving in complex real-world scenarios. In a separate prediction task, DriveGPT outperforms state-of-the-art baselines and exhibits improved performance by pretraining on a large-scale dataset, further validating the benefits of data scaling.
@@ 1 Introduction
While scaling up model and dataset sizes has been critical for recent advances in sequential modeling for text prediction [11], it remains unclear whether these scaling trends can be directly extended to behavior modeling, particularly in driving tasks, due to several unique challenges. First, driving tasks involve a wider range of input modalities, including agent trajectories and map information, unlike language tasks that rely solely on textual inputs. Second, behavior modeling demands spatial reasoning and an understanding of physical kinematics. Such capabilities are typically beyond the scope of language models. Finally, the collection of large-scale driving datasets requires substantial effort and resources, making it far more challenging than gathering textual data. As a result, existing work is often constrained by the availability of training data or the scalability of the models, as summarized in Table 1.
In our work, we present a comprehensive study of scaling up data sizes and model parameters in the context of behavior modeling for autonomous driving, which predicts future actions of traffic agents to support critical tasks such as planning and motion prediction. Specifically, we train a transformer-based autoregressive behavior model on over 100 million high-quality human driving examples, $\sim$50 times more than existing open-source datasets, and scale the model over 1 billion parameters, outsizing existing published behavior models.
@@
    We determine empirical driving scaling laws for an autoregressive behavior model in terms of data size, model parameters, and compute. We validate the value of scaling up training data and compute, and observe better model scalability as training data increases, consistent with the language scaling literature.
@@
    We quantitatively and qualitatively compare models from our scaling experiments to validate their effectiveness in real-world driving scenarios. We present real deployment of our model through closed-loop driving in challenging conditions.
@@ 2.1 Behavior Modeling
Behavior modeling is a critical task in autonomous driving, which covers a broad spectrum of tasks including planning, prediction, and simulation. Taking multimodal inputs including agent history states and map information, behavior models predict the future states of these traffic agents by reasoning about agent dynamics [2, 23], interactions [24, 9], and driving environments [12, 14].
Among learning-based models, transformers have gained popularity due to their ability to fuse multimodal inputs as encoders [17, 29, 31, 8] and model long-range temporal relationships as decoders [21, 22]. Despite the success of transformers in behavior modeling, existing literature is often restricted by the size of model parameters due to limited training data, which fails to capture the full scaling potential of transformer-based models. In our work, we scale up our transformer models to include billions of parameters, by training on a large-scale dataset including more than 100M driving demonstrations, and validate the scalability of transformer-based models in the context of autonomous driving.
@@ 2.2 Large Transformer Models
Recent work has studied the scalability of behavior models in the context of motion prediction [21], planning [25], and simulation [7], yet these studies are either constrained by limited data size (up to a couple of million training examples), or focused on a few orders of magnitude in terms of data and model scaling, limiting the potential to draw statistically significant conclusions over a large scaling range.
In this paper, we study the scaling properties (in terms of data samples, model parameters, and compute) across a much larger range compared to prior work. More specifically, we target an autoregressive decoder architecture that has been proven to be both scalable (as in the LLM literature) and effective in generating accurate trajectories for traffic agents [21].
Beyond autonomous driving, there is limited relevant literature on building large transformer models for robotic tasks [18, 26], which share a transformer architecture similar to our work. Robotics models often share different input features and dynamic models, and operate in different environments, making them difficult to apply directly to driving tasks.
@@ 3 Behavior Model
We use a standard encoder-decoder architecture as our behavior model, as shown in Fig. 2. We use transformer-based models as our encoder and decoder backbones due to their scalability in related sequential modeling tasks.
@@ 3.3 LLM-Style Trajectory Decoder
where $a_{t}$ is the predicted Verlet action, and $(s_{t}-s_{t-1})$ assumes a constant velocity step. This representation helps predict smooth trajectories using a small set of actions.
@@ 3.4 Training
We use a single cross-entropy classification loss over the action space, where the target action is selected as the one that is closest to the ground truth future trajectories. We refer the reader to Appendix A for more training details.
@@ 4 Scaling Experiments
The goal of our scaling experiments is to determine the effect of data and model size on behavior prediction performance. Quantifying scaling laws similar to those seen in LLMs [11] can help prioritize the value of data and compute for future research directions in behavior modeling. We focus our effort on exploring the next frontier of data and model size — over an order of magnitude beyond previously published work.
4.1 Large-Scale Driving Dataset
From millions of miles of high-quality real-world human driving demonstrations, we curated a small subset of 120M segments for an internal research dataset. The dataset is carefully curated and balanced to cover diverse geographic regions from multiple cities and countries, and a variety of challenging driving scenarios, such as passing a double-parked vehicle, unprotected left turns, construction zones, etc. We extracted map information, target agent states, and nearby agent states into vectorized representation, as customary in behavior modeling literature [4].
@@ 4.2 Data Scaling
Figure 3: Performance increases with dataset size across a range of model parameters, indicating that data is a limiting factor. Both axes are on a log scale. An exponential fit was applied to all data points except for the 1.5M curve, resulting in the following relationship: $\log(L)=-0.102\log(D)+2.663$ with an $R^{2}$ value of 0.986, where $L$ is the validation loss and $D$ is the number of unique training samples.
Data scaling results are summarized in Fig. 3. The smallest dataset of 2.2M samples mimics the size of Waymo Open Motion Dataset (WOMD) [3], a large open-source dataset for behavior modeling ($\sim$44k scenarios with multiple target agents per scenario). We select a few subsets of our internal research dataset to study data scaling across different orders of dataset sizes. Our experiments use $\sim$50x more data than WOMD, exploring a new region of the design space.
The results indicate that as the model is trained on more unique data samples, the performance improves, regardless of model size. Extrapolating from the scaling law in Fig. 3, to improve the best loss by another $10\%$, we would need to include $350$M more training examples. A $20\%$ improvement would require about $1.4$B more examples. As a result, we find that data remains the bottleneck for further improving driving performance.
@@ 4.3 Model Scaling
We now study model sizes across three orders of magnitude (1.5M to 1.4B parameters). We increase model size by increasing the hidden dimensions of transformers for simplicity. We notice that modifying other parameters such as number of attention heads and hidden dimensions per head does not lead to noticeable changes in the results, as studied in Appendix B.
Training larger models is sensitive to learning rates, as observed in other scaling studies [11, 6]. For each model size, we run multiple experiments at different learning rates to select the one with the optimal performance, as summarized in Appendix B.
Figure 4: Model scaling is more effective as training data increases. The validation loss improves up to $\sim$100M parameters when trained on the full dataset.
Results in Fig. 4 demonstrate that increasing the amount of training data enhances the effectiveness of model scaling. Specifically, when the dataset size is up to 21M, the impact on the validation loss is barely noticeable across a large range of model sizes. Beyond 21M samples, validation losses improve with larger models – up to 12M parameters when trained on the 42M dataset and up to 94M parameters when trained on the 120M dataset – before reaching a plateau and eventually overfitting. These findings further reinforce that data is the primary bottleneck for scaling models, aligning with observations in the LLM scaling literature [11].
@@ 4.4 Compute Scaling
Next, we investigate if there is an optimal combination of data size and model parameters for a fixed compute budget. Given the large computational expense for training models at the scales we are exploring, it is important to make the best use of our data. In this study, we fixed the compute budget at different FLOP groups. For a fixed compute budget, we can allocate resources either to model parameters or data samples, keeping their product constant.
Fig. 6 plots performance with different compute budgets. The trend clearly shows that a larger compute budget leads to better performance, with optimal model size increasing accordingly, as indicated by the “best” gray line. The results further reveal that data is the main bottleneck, as the smallest model outperforms others in the three largest FLOP groups.
@@ 5 Planning and Prediction Experiments
5.1 Internal Evaluation: AV Planning
We measure the planning performance on a comprehensive test set through a set of standard geometric metrics including minADE (mADE), minFDE (mFDE), and miss rate (MR). Additionally, we use semantic-based metrics including offroad rate (Offroad) that measures the ratio of trajectories that leave the road and collision rate (Collision) that measures the ratio of trajectories overlapping with static agents. We normalize these metrics across experiments to highlight relative performance changes.
@@ 5.1.1 Data Scaling Results
We compare a DriveGPT model at 26M parameters trained on datasets of different sizes. The baseline dataset (2.2M) is selected to mimic the size of a typical behavior modeling dataset such as WOMD.
The results are presented in Table 2, where we see that training on more data samples significantly improves the quality of predicted AV trajectories, in terms of critical semantics metrics in driving including offroad rate and collision rate, as well as geometric metrics. These improvements are consistent with Sec. 4.2.
@@ 5.1.2 Model Scaling Results
The results are presented in Table 6, where all metrics improve as the model size increases up to 94M parameters, with further gains in the collision metric at 163M parameters. Although validation loss plateaus at 94M in Fig. 4, driving metrics continue to improve with increased model capacities, highlighting the potential benefits of introducing more parameters for enhanced driving performance.
@@ 5.1.3 Closed-Loop Driving
We demonstrate the effectiveness of DriveGPT as a real-time motion planner deployed in a closed-loop setting. The model takes input features from an industry-level perception system that outputs agent states and map information.
In Fig. 10, we present a challenging example in dense urban traffic where there are two double-parked vehicles blocking the path forward along with other oncoming vehicles.  DriveGPT generates smooth and safe trajectories, bypassing the blocking vehicles and moving back to the original lane afterward. More examples are presented in Appendix C.1.
@@ 5.2 External Evaluation: Motion Prediction
HDGT [8] 0.7676 1.1077 0.1325 0.3709
MTR [22] 0.6050 1.2207 0.1351 0.4216
HPTR [29] 0.5565 1.1393 0.1434 0.3904
Wayformer Ensemble [17] 0.5454 1.1280 0.1228 0.4335
MotionLM Ensemble [21] 0.5509 1.1199 0.1058 0.4507
DriveGPT-WOMD 0.5279 1.0609 0.1236 0.3795
DriveGPT-Finetune 0.5240 1.0538 0.1202 0.3857
Table 4: On the WOMD test set,  DriveGPT achieves better results in geometric metrics without any ensembling, and improved performance after pretraining. Results are averaged over three agent types (vehicle, pedestrian, cyclist) and three prediction horizons (3s, 5s, 8s). The best metric is highlighted in bold and the second best is underlined.
@@
MTR 0.7642 1.5257 0.1514
DriveGPT-WOMD 0.6396 1.2823 0.1213
DriveGPT-Finetune 0.6326 1.2636 0.1181
@@
MTR 0.7022 1.4093 0.1786
DriveGPT-WOMD 0.6317 1.2617 0.1756
DriveGPT-Finetune 0.6294 1.2641 0.1691
[…]
MTR 0.3486 0.7270 0.0753
DriveGPT-WOMD 0.3122 0.6388 0.0738
DriveGPT-Finetune 0.3100 0.6338 0.0733
Table 5:  DriveGPT achieves better geometric metrics across a diverse set of agent types. Results are averaged over 3 prediction horizons.
@@ 5.2.1 Open-Source Encoder
For our external evaluation, we use the open-source MTR [22] encoder. This encoder is similar to the one described in Sec. 3.2. We make this change to improve the reproducibility of our results and take advantage of MTR’s open-source dataloading code for WOMD. We use the same autoregressive decoder as described in Sec. 3.3.
@@ 5.2.2 Pretraining Setup
We pretrain DriveGPT by training on our internal research dataset for one epoch (as in Sec. 4). We load the pretrained checkpoint and finetune the model using the same training setup as in the MTR codebase, where we train the model for 30 epochs using a weighted decay learning rate scheduler.
5.2.3 Results
@@
We measure model performance via a set of standard WOMD metrics, including minADE, minFDE, miss rate, and soft mAP. Each metric is measured on the test set and computed over three different time horizons.
@@
We present two variants of our method to validate its effectiveness on the motion prediction task, including DriveGPT-WOMD that is trained on WOMD, and DriveGPT-Finetune that is pretrained on our 120M internal research dataset and finetuned on WOMD. For baselines, we use a set of representative state-of-the-art models reported in [21].
We report results on the WOMD test set in Table 4. The results demonstrate that our method outperforms existing state-of-the-art non-ensemble models in terms of geometric metrics. Compared to Wayformer [17] and MotionLM [21] that use ensembles of up to 8 replicas, our model achieves the best minADE and minFDE metrics and the second-best miss rate metric without any ensembling.
@@
While we prioritize geometric metrics that emphasize the recall of predicted trajectory samples (i.e. not missing the critical trajectory), our model shows lower soft mAP scores due to suboptimal probability estimates. These estimates suffer from accumulated noises over time, as they are computed by compounding the action probabilities over a long sequence (as described in Eq. (1)). Consequently, we notice that the gap in soft mAP grows as the prediction horizon increases. This reveals a limitation of using an autoregressive decoder for accurate probability estimation, as also noted in the LLM literature [10, 5]. We defer improving the probability estimates of autoregressive models for behavior modeling as future work.
@@
We observe up to 3% additional gains by pretraining on our internal dataset, despite a large distribution shift between our internal dataset and the public WOMD dataset, in terms of trajectory distributions, feature noises, and differences in semantic definitions.
Diving deeper into agent-specific results in Table 5, we see consistent improvements across all agent types, compared to MTR that shares the same encoder as ours. This further validates the generalizability of our method, in addition to vehicle behavior modeling results described in Sec. 5.1.
5.2.4 Qualitative Comparison
6 Conclusion
We introduced DriveGPT, an LLM-style autoregressive behavior model, to better understand the effects of model parameters and dataset size for autonomous driving. We systematically explored model performance as a function of both dataset size and model parameters, uncovering LLM-like scaling laws for data and compute scaling, along with limited model scaling properties. We showed the quantitative and qualitative benefits of scaling for planning in real-world driving scenarios. Additionally, we demonstrated our method on a public motion prediction benchmark, where DriveGPT outperformed a state-of-the-art baseline by a large margin and achieved improved performance through pretraining on a large-scale dataset.
Appendix A Training Detail
A.1 Internal Evaluation: AV Planning
Our models follow the implementations described in [17, 21], and are trained using a batch size of 2048 and a standard Adam optimizer adopted in [11]. We follow the optimal learning rate schedule discovered in [6], which applies a cosine decay with a cycle length equivalent to the total number of training steps. All models were trained on 16 NVIDIA H100 GPUs.
A.2 External Evaluation: Motion Prediction
Appendix B Scaling Detail
B.1 Scaling Configurations
B.2 Ablation Study on Attention Heads
Model Size Hidden Dimension ($d_{\text{model}}$) Max LR
1.5M 64 0.0050
4M 128 0.0020
8M 192 0.0014
12M 256 0.0010
26M 384 0.0010
94M 768 0.0007
163M 1024 0.0004
358M 1536 0.0002
629M 2048 0.0001
1.4B 3072 0.0001
Figure 13: Scaling only the decoder exhibits a similar trend but worse results compared to scaling both encoder and decoder. Scaling only the encoder has little impact beyond 10M parameters.
B.3 Ablation Study on Decoder Scaling
In Fig. 13, we present model scaling results (in dashed lines) where we only scale up the autoregressive decoder, while keeping the encoder at a fixed size. Compared to scaling both encoder and decoder (in solid lines), scaling only the decoder exhibits a similar trend but yields worse performance, especially for models exceeding 10M parameters in two larger datasets. Therefore, our main results focus on scaling both encoder and decoder.
B.4 Ablation Study on Encoder Scaling
In Fig. 13, we present model scaling results (in dotted lines) where we only scale up the encoder, while keeping the decoder at a fixed size. The results indicate that encoder scaling is not as effective as decoder scaling.
B.5 Scaling From a Perplexity View
Figure 15: We fit the best perplexity values from all models in Fig. 3 as a function of dataset size. A simple linear regression yields $\log(P)=-0.28\log(D)+7.22$ with an $R^{2}$ value of 0.994, where $P$ is the validation perplexity and $D$ is the number of unique training examples.
In Fig. 14, we re-plot the scaling results, adding a secondary $y$-axis to represent perplexity. In Fig. 15, we plot the perplexity of our best models as a function of unique training examples. Extrapolating from the perplexity fit in the figure, a perplexity of 7 would require about $70$M more training examples and a perplexity of 5 would require $500$M more training examples.
Appendix C Additional Qualitative Examples
C.1 Additional Closed-Loop Driving Examples
C.2 Failure Cases on WOMD
Appendix D Additional WOMD Validation Results
In this section, we present additional ablation studies on the WOMD validation set, due to the submission limit of the WOMD test server.
D.1 WOMD Data Scaling in Pretraining